# Age-Gender-Country-Specific Death Rates Modelling and Forecasting: A Linear Mixed-Effects Model

### v1

@@ 3 The ASDRs Model for Multi Populations
In this section, we give an approach to modelling and forecasting ASDRs in an international context instead of looking at one country at a time. In the next section, we will consider ASDRs of a single population as a particular case of this section. We will consider the six countries chosen by Enchev et al. (2017). The population sizes of these countries are similar, their geographical locations are close, and there has been a remarkable drop in human mortality rates in the six countries since 1961. In Figure 4, vertical violin plots display the distribution of log mortality rates across the six European countries. Wider sections indicate a higher likelihood of observing values within that range, while narrower sections suggest a lower likelihood. The plots show remarkable similarity, indicating a shared distribution pattern for log mortality rates among the countries. The consistent lengths of the violin plots reflect the consistent frequency of observations for specific log mortality values across all countries. Notably, the Denmark plot exhibits a taller height, indicating greater variability in log mortality rates compared to the other countries. This implies a wider range of small mortality rate values in Denmark.
@@
We will use ASDRs for conventional 5-year age groups up to $110^{+}$ by single calendar years $1961-2019$. We will group individuals’ mortality rates by age, gender, and country.
@@
where $c\in\{{\rm AUT},{\rm BEL},{\rm CHE},{\rm CZE},{\rm DNK},{\rm SWE}\}$, $g={\rm Female},{\rm Male}$, $x=0,1,5$$\cdots$, $110^{+}$, and $t=1961,1962,\cdots,2010$. Because of the quadratic trend that we can see in Figure 5 it is the model (42) that is conceptually most appropriate for the dataset. The mixed-effects model combines a model for the fixed effects and a model for the random effects.
@@ 3.1 A Candidate Model Set
In this paper, females serve as the default gender, and age group 00 is established as the reference point for comparison. Consequently, the parameter estimates derived for the interaction of gender and age, and age groups in the fixed effects of the intercept of the model are intended to quantify the disparities between males and females and evaluate the differences between each age group $1$, $5$$\cdots$, and $110$ and age group 00.
@@ 3.2 Addressing Heteroskedasticity and Non-Linear QQ Plots in Mortality Models
After fitting the candidate model, we conducted a thorough examination of the residuals to assess the presence of heteroskedasticity, which refers to the unequal variance of residuals across a range of observed values (see James et al., 2013, chap. 3). Heteroskedasticity can manifest as a non-uniform scatter of residuals, which can impact the reliability of the model. Our analysis revealed that the residuals exhibited non-normality and heteroskedasticity, indicating that the variance of residuals was not consistent across different predicted values. To address this issue, we employed a data-cleaning procedure by excluding data points with absolute residuals greater than 0.1, signifying instances where the predicted and actual values diverged significantly. Subsequently, we re-estimated the LME model using this refined dataset, which typically retained about $85\%$ of the original data. This procedure resulted in improved characteristics of the residuals. We can observe the plot of the fitted values versus residuals and the QQ plot of residuals before and after these corrections in Figure 6. All findings presented in this paper are derived from the LME model post-critique, following the exclusion of data points and model refitting.
@@
It is worth noting that these two issues—heteroskedasticity and nonlinearity of the QQ plot (Wilk and Gnanadesikan, 1968; Thode, 2002) of the residuals—are observed not only in the candidate model but also in other mortality models such as the LC model, LL model, and the Plat model. We conducted checks for these issues in all three mortality models. We can see the manifestation of these problems for the LC model in Figure 7.
@@
We can also explore the application of a robust linear mixed-effects model using the robustlmm package (Koller, 2016) as a robust approach to addressing heteroskedasticity in our dataset. Unlike traditional linear modeling methods, robustlmm is designed to handle situations where the variability of the data points (residuals) is not consistent, which is a common issue in many real-world datasets. By incorporating robustlmm, we can effectively account for and mitigate the impact of this non-constant variance in the residuals. Importantly, this approach eliminates the need to eliminate or exclude specific data points from our dataset.
@@ 3.3 Stepwise Model Selection
We will use the refined dataset from $1961$ to $2010$ as the training set to estimate the parameters of the candidate model set. The REML estimation technique is used to fit the model (43) to the dataset. To do so, we will use the lmer function in the lme4 package in R Programming language (Bates et al., 2015). How can we determine which terms in the model (43) are necessary to fit our data better? Two goodness-of-fit measures that we can use are Akaike Information Criterion (AIC) (Sakamoto et al., 1986) and Schwarz’s Bayesian Information Criterion (BIC) (Schwarz, 1978). The following formulas give AIC and BIC:
@@
where $d$ is the total number of parameters, and $N$ is the sample size of the training set. To keep the model (43) maximal and remove further terms based on AIC/BIC, one algorithm we can use is known as stepwise selection. Comparing a restricted set of models, it picks the best model (Hastie et al., 2009). We implement a backward stepwise selection strategy that considers backward moves at each step and selects the best model (see Kuznetsova et al., 2017). According to the backward stepwise selection, the best-fit model for the log ASDRs of the twelve populations is the model
@@
One of the notable strengths of the LME model is its capability for stepwise model selection, allowing us to identify which terms are essential for improving data fit. However, it is important to keep in mind that the random effects are designed with a mean of 0, which means that any non-zero mean for a term in the random effects must be integrated into the fixed-effects terms (see Pinheiro and Bates, 2006, chap. 2). Consequently, we may need to consider incorporating new covariates as fixed effects if warranted.
@@ 3.4 Estimation of Model Parameters
$x$ $\hat{\beta}_{0x}$ sd$(\hat{\beta}_{0x})$ $\hat{\beta}_{0Mx}$ sd$(\hat{\beta}_{0Mx})$ $\hat{\beta}_{1Fx}$ sd$(\hat{\beta}_{1Fx})$ $\hat{\beta}_{1Mx}$ sd$(\hat{\beta}_{1Mx})$ $\hat{\beta}_{2Fx}$ sd$(\hat{\beta}_{2Fx})$ $\hat{\beta}_{2Mx}$ sd$(\hat{\beta}_{2Mx})$
@@
$\begin{split}\hat{\Psi}&=\begin{bmatrix}\quad 4.408*10^{-3}&-3.424*10^{-4}&-8.856*10^{-7}\\ -3.424*10^{-4}&\quad 1.264*10^{-3}&-1.344*10^{-5}\\ -8.856*10^{-7}&-1.344*10^{-5}&\quad 1.476*10^{-7}\\ \end{bmatrix}.\end{split}$ (49)
@@
The histograms displayed in Figure 9 illustrate normal distributions with mean values that closely approximate zero for the intercept, $k_{t}^{2}$, and cohort random effects. Specifically, the means for these effects are approximately $3.76*10^{-12}$, $-1.96*10^{-11}$, and $2.10*10^{-13}$, respectively.
@@
Table 2 presents a comprehensive overview of the ANOVA analysis conducted to evaluate the significance of the fixed terms within the LME model (see Kuznetsova et al., 2017). The p-values associated with all six fixed terms indicate their statistical significance.
Fixed-Effect Term Sum Sq Mean Sq NumDF DenDF F value Pr($>$F)
$x$ 4.34 0.19 23.00 10710.71 105.74 $<$ 2.2e-16 ******* * *
$g:x$ 2.11 0.09 24.00 10698.14 49.28 $<$ 2.2e-16 ******* * *
$g:x:{\rm{I}}(k_{ct})$ 11.91 0.25 48.00 $9$.26 138.91 $<$ 2.2e-16 ******* * *
${\rm{I}}(k_{t}^{2})$ 0.01 0.01 1.00 593.35 3.95 0.0473 ***
$g:x:{\rm{I}}(k_{ct}^{2})$ 10.83 0.23 48.00 8825.68 126.34 $<$ 2.2e-16 ******* * *
$\rm{cohrot}$ 0.04 0.04 1.00 11084.33 19.60 $<$ 2.2e-16 ******* * *
@@
Figures 12 and 13 display the LME and LC fits for log ASDRs of the twelve-population dataset, respectively, for both the training and test sets covering the years $1961-2019$. The models effectively capture the mortality rate patterns associated with age, gender, and country, indicating their overall adequacy. Remarkably, the LME model outperforms the LC model in accurately modelling infant mortality rates. The LME model demonstrates a better ability to capture the rates and patterns of infant mortality, resulting in more precise predictions of life expectancies, as we will see in section 3.6.
Figure 12: LME fits and predictions of log ASDRs for the twelve-population dataset ($1961-2019$).
Figure 13: LC fits and predictions of log ASDRs for the twelve-population dataset ($1961-2019$).
@@ 3.5 The ASDRs Model Forecasts
We will use the data of $2011-2019$ to assess the predictive quality of the model (46) that the stepwise algorithm has selected. We have found that the time-dependent covariates $k_{ct}$ and $k_{t}$ can be modeled and forecast using ARIMA models (see Box et al., 2015). We assume random walks with drift (i.e., ARIMA(0,1,0)) for $k_{ct}$ and $k_{t}$:
@@
Figure 16: Observed and predicted values of $k_{ct}$ ($1961-2019$).
Figure 17: Observed and predicted values of $k_{t}$ ($1961-2019$).
Log ASDRs forecasts are generated using estimated parameters: $\beta_{0}$, $\beta_{0x}$, $\beta_{0gx}$, $\beta_{1gx}$, $\beta_{2}$, $\beta_{2gx}$, $\gamma_{{\rm{cohort}}}$, $\eta_{0cgx}$, $\eta_{2cgx}$, and $\eta_{{\rm{cohort}}_{cgx}}$. These parameters, along with the forecasted values for $k_{ct}$ and $k_{t}$, are utilized in the forecasting process. It is assumed that these parameters remain constant and do not change over time. Actuarial applications have shown the importance of estimating uncertainty in mortality forecasting, see Olivieri (2001); Li et al. (2009).
To predict mortality rates using the LME model for the year 2019 and quantify their uncertainty, we employ a simulation approach utilizing the predictInterval function in the R programming language’s merTools package (Knowles et al., 2023). We calculate prediction intervals for 2019, which incorporate the uncertainty in fixed coefficients, random effects, and residual variance, with $95\%$ coverage based on 1,000 simulations, and we use the median as the point estimate.
An alternative approach is to perform a bootstrap (see Davison and Hinkley, 1997; Efron, 1994) simulation utilizing the bootMer function from the R programming language’s lme4 package (see Bates et al., 2015). This method allows for the estimation of prediction intervals by resampling and is considered a gold-standard approach for capturing the uncertainty inherent in LMMs. However, it is worth noting that obtaining a sufficient number of replications from bootMer can be computationally time-consuming, especially when the initial model fit requires a substantial amount of time, such as hours instead of seconds.
@@
As a measure of predictive accuracy, mean square error (MSE) is discussed. The MSE of the twelve populations for the test set ($2011-2019$) are shown in Table 3. We have fitted the LC model to each population separately, and we will use the demographic LC model as a benchmark model to evaluate the performance of the LME model. In the test datasets, the MSEs of the LME model are lower than those of the LC models for 11 out of the 12 populations.
@@ 3.8 The All-in-One LME Model: Integrating Age, Gender, Country, Cohort, and GDP per Capita Effects
Boonen and Li (2017) have introduced a novel approach to mortality modelling for multiple populations. Instead of relying solely on historical mortality data and latent factors, the authors propose a multipopulation stochastic mortality model that incorporates the influence of economic growth, specifically using real GDP per capita. Building upon the LL model, this extended framework aims to capture common mortality trends among populations with similar socioeconomic conditions. The incorporation of GDP into mortality models is of paramount importance as it recognizes the substantial impact of economic factors on mortality patterns, providing a more comprehensive understanding of population dynamics and enhancing the accuracy of mortality projections.
@@
To fit the model using the lmer function in the lme4 package in the R programming language (see Bates et al., 2015), the LME model can be specified as follows:
@@ 4 The ASDRs Model for Single Population
Let $m_{x,t}$ denote the ASDR at age $x$ and time $t$, for $x=0,1,\cdots,\omega$ and $t=0,1,\cdots,T$, see Dickson et al. (2019). Let $y_{xt}=\log(m_{x,t})$, the log of the observed age-specific death rates in a given year $t$. Let $k_{2_{t}}$ represent the average of $y_{xt}$ for two distinct age range groups: the average for age groups 0 to 40 and the average for age groups 41 to $\omega$ at time $t$. Additionally, let the average of $y_{xt}$ at time $t$ be $k_{t}$:
@@
$y_{x}$ $=\begin{bmatrix}y_{x0}\\ y_{x1}\\ \vdots\\ y_{xT}\end{bmatrix},$ (59)
$X_{x}$ $=\begin{bmatrix}1&k_{0}&k_{2_{0}}&\cdots&k_{0}^{n_{1}}&k_{2_{0}}^{n_{1}}&0-x\\ 1&k_{1}&k_{2_{1}}&\cdots&k_{1}^{n_{1}}&k_{2_{1}}^{n_{1}}&1-x\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ 1&k_{T}&k_{2_{T}}&\cdots&k_{T}^{n_{1}}&k_{2_{T}}^{n_{1}}&T-x\end{bmatrix},$ $\beta$ $=\begin{bmatrix}\beta_{0}\\ \beta_{1}\\ \beta_{2_{1}}\\ \vdots\\ \beta_{{n_{1}}}\\ \beta_{2_{n_{1}}}\\ \gamma_{\rm{cohort}}\end{bmatrix},$ (71)
$Z_{x}$ $=\begin{bmatrix}1&k_{0}&k_{0}^{2}&\cdots&k_{0}^{n_{2}}&0-x\\ 1&k_{1}&k_{1}^{2}&\cdots&k_{1}^{n_{2}}&1-x\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 1&k_{T}&k_{T}^{2}&\cdots&k_{T}^{n_{2}}&T-x\end{bmatrix},$ $\eta_{x}$ $=\begin{bmatrix}\eta_{0x}\\ \eta_{1x}\\ \vdots\\ \eta_{n_{2}x}\\ \eta_{{\rm{cohort}}_{x}}\end{bmatrix},$ (81)
@@
where $n_{1}$ and $n_{2}$ are non-negative integers; $\Psi$ is a positive-definite, symmetric $(n_{2}+2)\times(n_{2}+2)$ matrix; $\beta$ is the fixed effects vector, and $\eta_{x}$ is the random effects vector; $X_{x}$ and $Z_{x}$ are known regressor matrices for the fixed-effects and random-effects; and $\epsilon_{x}$ is the within-age-group error vector. We assume that the random effects $\eta_{x}$ and the within-age-group errors $\epsilon_{x}$ are independent for different age groups. We also assume they are independent of one another for the same age group (see Pinheiro and Bates, 2006, chap. 2).
@@
for further details, see Wang (2013).
@@
The REML estimation technique is used to fit the model (87) to each individual dataset (see Bates et al., 2015) .
@@
Figure 26: Ratio visualization of LC MSE divided by LME MSE (log ASDRs). Ratios greater than $1$ signify that the LC MSE is larger than the LME MSE, implying potential higher prediction errors in the LC model relative to the LME model.

### v2

@@ 3 The ASDRs Model for Multi Populations
In this section, we give an approach to modelling and forecasting ASDRs in an international context instead of looking at one country at a time. In the next section, we will consider ASDRs of a single population as a particular case of this section. We will consider the six countries chosen by enchev2017multi. The population sizes of these countries are similar, their geographical locations are close, and there has been a remarkable drop in human mortality rates in the six countries since 1961. In Figure 4, vertical violin plots display the distribution of log mortality rates across the six European countries. Wider sections indicate a higher likelihood of observing values within that range, while narrower sections suggest a lower likelihood. The plots show remarkable similarity, indicating a shared distribution pattern for log mortality rates among the countries. The consistent lengths of the violin plots reflect the consistent frequency of observations for specific log mortality values across all countries. Notably, the Denmark plot exhibits a taller height, indicating greater variability in log mortality rates compared to the other countries. This implies a wider range of small mortality rate values in Denmark.
@@
We will use ASDRs for conventional 5-year age groups up to $110^{+}$ by single calendar years 196120191961-2019. We will group individuals’ mortality rates by age, gender, and country.
@@
[…]
@@ 3.1 A Candidate Model Set
[…]
@@ 3.2 Addressing Heteroskedasticity and Non-Linear QQ Plots in Mortality Models
After fitting the candidate model, we conducted a thorough examination of the residuals to assess the presence of heteroskedasticity, which refers to the unequal variance of residuals across a range of observed values (see james2013introduction, chap. 3). Heteroskedasticity can manifest as a non-uniform scatter of residuals, which can impact the reliability of the model. Our analysis revealed that the residuals exhibited non-normality and heteroskedasticity, indicating that the variance of residuals was not consistent across different predicted values. To address this issue, we employed a data-cleaning procedure by excluding data points with absolute residuals greater than 0.1, signifying instances where the predicted and actual values diverged significantly. Subsequently, we re-estimated the LME model using this refined dataset, which typically retained about $85\%$ of the original data. This procedure resulted in improved characteristics of the residuals. We can observe the plot of the fitted values versus residuals and the QQ plot of residuals before and after these corrections in Figure 6. All findings presented in this paper are derived from the LME model post-critique, following the exclusion of data points and model refitting.
@@
It is worth noting that these two issues—heteroskedasticity and nonlinearity of the QQ plot (wilk1968probability; thode2002testing) of the residuals—are observed not only in the candidate model but also in other mortality models such as the LC model, LL model, and the Plat model. We conducted checks for these issues in all three mortality models. We can see the manifestation of these problems for the LC model in Figure 7.
@@
We can also explore the application of a robust linear mixed-effects model using the robustlmm package (koller2016robustlmm) as a robust approach to addressing heteroskedasticity in our dataset. Unlike traditional linear modeling methods, robustlmm is designed to handle situations where the variability of the data points (residuals) is not consistent, which is a common issue in many real-world datasets. By incorporating robustlmm, we can effectively account for and mitigate the impact of this non-constant variance in the residuals. Importantly, this approach eliminates the need to eliminate or exclude specific data points from our dataset.
@@ 3.3 Stepwise Model Selection
We will use the refined dataset from $1961$ to $2010$ as the training set to estimate the parameters of the candidate model set. The REML estimation technique is used to fit the model (10) to the dataset. To do so, we will use the lmer function in the lme4 package in R Programming language (JSSv067i01). How can we determine which terms in the model (10) are necessary to fit our data better? Two goodness-of-fit measures that we can use are Akaike Information Criterion (AIC) (sakamoto1986akaike) and Schwarz’s Bayesian Information Criterion (BIC) (schwarz1978estimating). The following formulas give AIC and BIC:
@@
where $d$ is the total number of parameters, and $N$ is the sample size of the training set. To keep the model (10) maximal and remove further terms based on AIC/BIC, one algorithm we can use is known as stepwise selection. Comparing a restricted set of models, it picks the best model (hastie2009elements). We implement a backward stepwise selection strategy that considers backward moves at each step and selects the best model (see kuznetsova2017lmertest). According to the backward stepwise selection, the best-fit model for the log ASDRs of the twelve populations is the model
@@
One of the notable strengths of the LME model is its capability for stepwise model selection, allowing us to identify which terms are essential for improving data fit. However, it is important to keep in mind that the random effects are designed with a mean of 0, which means that any non-zero mean for a term in the random effects must be integrated into the fixed-effects terms (see pinheiro2006mixed, chap. 2). Consequently, we may need to consider incorporating new covariates as fixed effects if warranted.
@@ 3.4 Estimation of Model Parameters
[…]
@@
[…]
@@
The histograms displayed in Figure 9 illustrate normal distributions with mean values that closely approximate zero for the intercept, $k_{t}^{2}$, and cohort random effects. Specifically, the means for these effects are approximately $3.76*10^{-12}$, $1.96*10^{-11}$, and $2.10*10^{-13}$, respectively.
@@
Table 2 presents a comprehensive overview of the ANOVA analysis conducted to evaluate the significance of the fixed terms within the LME model (see kuznetsova2017lmertest). The p-values associated with all six fixed terms indicate their statistical significance.
Fixed-Effect Term Sum Sq Mean Sq NumDF DenDF F value Pr(>>F)
$x$ 4.34 0.19 23.00 10710.71 105.74 << 2.2e-16 ***
$g:x$ 2.11 0.09 24.00 10698.14 49.28 << 2.2e-16 ***
$g:x:{\rm{I}}(k_{ct})$ 11.91 0.25 48.00 9999.26 138.91 << 2.2e-16 ***
${\rm{I}}(k_{t}^{2})$ 0.01 0.01 1.00 593.35 3.95 0.0473 *
$g:x:{\rm{I}}(k_{ct}^{2})$ 10.83 0.23 48.00 8825.68 126.34 << 2.2e-16 ***
$\rm{cohrot}$ 0.04 0.04 1.00 11084.33 19.60 << 2.2e-16 ***
@@
Figures 12 and 13 display the LME and LC fits for log ASDRs of the twelve-population dataset, respectively, for both the training and test sets covering the years 196120191961-2019. The models effectively capture the mortality rate patterns associated with age, gender, and country, indicating their overall adequacy. Remarkably, the LME model outperforms the LC model in accurately modelling infant mortality rates. The LME model demonstrates a better ability to capture the rates and patterns of infant mortality, resulting in more precise predictions of life expectancies, as we will see in section 3.6.
Figure 12: LME fits and predictions of log ASDRs for the twelve-population dataset (196120191961-2019).
Figure 13: LC fits and predictions of log ASDRs for the twelve-population dataset (196120191961-2019).
@@ 3.5 The ASDRs Model Forecasts
We will use the data of 201120192011-2019 to assess the predictive quality of the model (13) that the stepwise algorithm has selected. We have found that the time-dependent covariates $k_{ct}$ and $k_{t}$ can be modeled and forecast using ARIMA models (see box2015time). We assume random walks with drift (i.e., ARIMA(0,1,0)) for $k_{ct}$ and $k_{t}$:
@@
Figure 16: Observed and predicted values of $k_{ct}$ (196120191961-2019).
Figure 17: Observed and predicted values of $k_{t}$ (196120191961-2019).
Log ASDRs forecasts are generated using estimated parameters: $\beta_{0}$, $\beta_{0x}$, $\beta_{0gx}$, $\beta_{1gx}$, $\beta_{2}$, $\beta_{2gx}$, $\gamma_{{\rm{cohort}}}$, $\eta_{0cgx}$, $\eta_{2cgx}$, and $\eta_{{\rm{cohort}}_{cgx}}$. These parameters, along with the forecasted values for $k_{ct}$ and $k_{t}$, are utilized in the forecasting process. It is assumed that these parameters remain constant and do not change over time. Actuarial applications have shown the importance of estimating uncertainty in mortality forecasting, see olivieri2001uncertainty; li2009uncertainty.
To predict mortality rates using the LME model for the year 2019 and quantify their uncertainty, we employ a simulation approach utilizing the predictInterval function in the R programming language’s merTools package (knowles2023package). We calculate prediction intervals for 2019, which incorporate the uncertainty in fixed coefficients, random effects, and residual variance, with $95\%$ coverage based on 1,000 simulations, and we use the median as the point estimate.
An alternative approach is to perform a bootstrap (see Davison1997Bootstrap; efron1994introduction) simulation utilizing the bootMer function from the R programming language’s lme4 package (see JSSv067i01). This method allows for the estimation of prediction intervals by resampling and is considered a gold-standard approach for capturing the uncertainty inherent in LMMs. However, it is worth noting that obtaining a sufficient number of replications from bootMer can be computationally time-consuming, especially when the initial model fit requires a substantial amount of time, such as hours instead of seconds.
@@
As a measure of predictive accuracy, mean square error (MSE) is discussed. The MSE of the twelve populations for the test set (201120192011-2019) are shown in Table 3. We have fitted the LC model to each population separately, and we will use the demographic LC model as a benchmark model to evaluate the performance of the LME model. In the test datasets, the MSEs of the LME model are lower than those of the LC models for 11 out of the 12 populations.
@@ 3.8 The All-in-One LME Model: Integrating Age, Gender, Country, Cohort, and GDP per Capita Effects
boonen2017modeling have introduced a novel approach to mortality modelling for multiple populations. Instead of relying solely on historical mortality data and latent factors, the authors propose a multipopulation stochastic mortality model that incorporates the influence of economic growth, specifically using real GDP per capita. Building upon the LL model, this extended framework aims to capture common mortality trends among populations with similar socioeconomic conditions. The incorporation of GDP into mortality models is of paramount importance as it recognizes the substantial impact of economic factors on mortality patterns, providing a more comprehensive understanding of population dynamics and enhancing the accuracy of mortality projections.
@@
To fit the model using the lmer function in the lme4 package in the R programming language (see JSSv067i01), the LME model can be specified as follows:
@@ 4 The ASDRs Model for Single Population
Let $m_{x,t}$ denote the ASDR at age $x$ and time $t$, for $x=0,1,\cdots,\omega$ and $t=0,1,\cdots,T$, see dickson2019actuarial. Let $y_{xt}=\log(m_{x,t})$, the log of the observed age-specific death rates in a given year $t$. Let $k_{2_{t}}$ represent the average of $y_{xt}$ for two distinct age range groups: the average for age groups 0 to 40 and the average for age groups 41 to $\omega$ at time $t$. Additionally, let the average of $y_{xt}$ at time $t$ be $k_{t}$:
@@
$y_{x}$ $=\begin{bmatrix}y_{x0}\\ y_{x1}\\ \vdots\\ y_{xT}\end{bmatrix}$,
$X_{x}$ $=\begin{bmatrix}1&k_{0}&k_{2_{0}}&\cdots&k_{0}^{n_{1}}&k_{2_{0}}^{n_{1}}&0-x\\ 1&k_{1}&k_{2_{1}}&\cdots&k_{1}^{n_{1}}&k_{2_{1}}^{n_{1}}&1-x\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ 1&k_{T}&k_{2_{T}}&\cdots&k_{T}^{n_{1}}&k_{2_{T}}^{n_{1}}&T-x\end{bmatrix}$, $\beta$ $=\begin{bmatrix}\beta_{0}\\ \beta_{1}\\ \beta_{2_{1}}\\ \vdots\\ \beta_{{n_{1}}}\\ \beta_{2_{n_{1}}}\\ \gamma_{\rm{cohort}}\end{bmatrix}$,
$Z_{x}$ $=\begin{bmatrix}1&k_{0}&k_{0}^{2}&\cdots&k_{0}^{n_{2}}&0-x\\ 1&k_{1}&k_{1}^{2}&\cdots&k_{1}^{n_{2}}&1-x\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 1&k_{T}&k_{T}^{2}&\cdots&k_{T}^{n_{2}}&T-x\end{bmatrix}$, $\eta_{x}$ $=\begin{bmatrix}\eta_{0x}\\ \eta_{1x}\\ \vdots\\ \eta_{n_{2}x}\\ \eta_{{\rm{cohort}}_{x}}\end{bmatrix}$,
@@
where $n_{1}$ and $n_{2}$ are non-negative integers; $\Psi$ is a positive-definite, symmetric $(n_{2}+2)\times(n_{2}+2)$ matrix; $\beta$ is the fixed effects vector, and $\eta_{x}$ is the random effects vector; $X_{x}$ and $Z_{x}$ are known regressor matrices for the fixed-effects and random-effects; and $\epsilon_{x}$ is the within-age-group error vector. We assume that the random effects $\eta_{x}$ and the within-age-group errors $\epsilon_{x}$ are independent for different age groups. We also assume they are independent of one another for the same age group (see pinheiro2006mixed, chap. 2).
@@
for further details, see wang2013identifiability.
@@
The REML estimation technique is used to fit the model (23) to each individual dataset (see JSSv067i01) .
@@
Figure 26: Ratio visualization of LC MSE divided by LME MSE (log ASDRs). Ratios greater than 11 signify that the LC MSE is larger than the LME MSE, implying potential higher prediction errors in the LC model relative to the LME model.